OpenStudy (anonymous):

Television sets. What does it mean to refer to a 20-in TV set or a 25-in TV set? Such units refer to the diagonal of the screen. A 35-in TV set also has a width of 28 inches. What is the height in inches, of the 35-in TV?

OpenStudy (anonymous):

you should use phythagoras: A^2+B^2=C^2 Givem A=28 (width) and C=35 (diagonal), find B (length)

OpenStudy (anonymous):

is the answer 441? I'm a bit confused.

OpenStudy (anonymous):

B^2 is 441. You need B

OpenStudy (anonymous):

21

OpenStudy (anonymous):

I squared 441. Was that the correct move?

OpenStudy (anonymous):

B=21 is the correct answer.

OpenStudy (anonymous):

Thank you :)

OpenStudy (anonymous):

you didn't square 441, you took 441's square root , which is 21 sqrt(441) = 21

OpenStudy (anonymous):

oops :)